# POP: Prefill-Only Pruning for Efficient Large Model Inference

### v1

@@ 3.3 Prefill-Only Pruning for Efficient Inference
A potential concern regarding independent KV projections is the representation mismatch for deep layers. We provide further experiments and discussions in Appendix C to address this concern.
@@ 4.3 Inference Speedup
Table 2: Prefill speedup comparison across different models and input lengths. All experiments are conducted with a batch size of 8. Values represent the speedup ratio relative to the full model ($1.0\times$).
@@ 4.4 Ablation Study
Effectiveness of Design Choices. We first verify the necessity of our three key design components: (1) targeting deep layers, (2) independent KV projections, and (3) boundary handling. Experimental results are shown in Table 3; detailed discussions are presented in Appendix B.
@@
Sensitivity to Pruning Ratio. We further investigate the trade-off between inference efficiency and model performance by varying the pruning ratio from 20% to 60%. The prefill speedup is measured with a sequence length of 1024 and a batch size of 4. Experimental Results are presented in Table 4.
@@ 7 Limitations
Second, our current implementation is based on a monolithic inference pipeline modified from the Transformers library. Recent advancements in inference systems propose disaggregated systems that deploy prefill and decode instances on separate hardware resources (Zhong et al., 2024; Patel et al., 2024). As POP naturally treats these two stages differently, it holds great potential for integration into these systems to further maximize cluster-level throughput. However, adapting POP for such distributed frameworks involves non-trivial engineering efforts, which we leave for future research.
Appendix A Details on Evaluation Strategies
Appendix B Effectiveness of Design Choices
To verify the necessity of our three key design components: (1) targeting deep layers, (2) independent KV projections, and (3) boundary handling, we conduct extensive experiments on Qwen3-VL. Experimental results are shown in Table 3.
Appendix C Robustness to Representation Mismatch

### v2

@@ 3.3 Prefill-Only Pruning for Efficient Inference
A potential concern regarding independent KV projections is the representation mismatch for deep layers. We provide further experiments and discussions in Appendix A to address this concern.
@@ 4.3 Inference Speedup
Table 2: TTFT speedup comparison across different models and input lengths. All experiments are conducted with a batch size of 8. Values represent the speedup ratio relative to the full model ($1.0\times$).
@@ 4.4 Ablation Study
4.4.1 Effectiveness of Design Choices
We first verify the necessity of our three key design components: (1) targeting deep layers, (2) independent KV projections, and (3) boundary handling. Experimental results are shown in Table 3.
@@ 4.4.1 Effectiveness of Design Choices
4.4.2 Sensitivity to Pruning Ratio
We further investigate the trade-off between inference efficiency and model performance by varying the pruning ratio from 20% to 60%. The prefill speedup is measured with a sequence length of 1024 and a batch size of 4. Experimental Results are presented in Table 4.
@@ 7 Limitations
Second, our current implementation is based on a monolithic inference pipeline modified from the Transformers library. Recent advancements in inference systems such as DistServe (Zhong et al., 2024) and Splitwise (Patel et al., 2024) propose disaggregated systems that deploy prefill and decode instances on separate hardware resources. As POP naturally treats these two stages differently, it holds great potential for integration into these systems to further maximize cluster-level throughput. However, adapting POP for such distributed frameworks involves non-trivial engineering efforts, which we leave for future research.
Acknowledgments
We thank all the reviewers for their insightful comments. This work is supported by the National Natural Science Foundation of China (No. 62472330).
Appendix A Robustness to Representation Mismatch
@@
Appendix B Integration with Orthogonal Methods
Table 5: Accuracies on long context tasks and TTFT speedup ratios. Experiments are conducted with greedy decoding (Temperature = 0). TTFT speedup is measured with batch size 1, input length 32K.
Method TTFT Speedup HotpotQA MultiFieldQA
[…]
As noted in our Related Work section (Section 5), POP (model pruning) is orthogonal to both sparse attention and token pruning methods. Technically, while token pruning reduces the sequence length $N$ (temporal dimension) and sparse attention reduces computation within attention blocks (computational sparsity), POP physically removes entire transformer layers $L$ (depth dimension). Thus, POP reduces the computational cost per token regardless of the sequence length or attention pattern, making it naturally complementary to these approaches.
To demonstrate this, we integrate POP with FlexPrefill Lai et al. (2025) and LLMLingua-2 Pan et al. (2024). Considering that both baselines are primarily designed to handle long-context challenges, we conducted additional experiments on Llama-3.1-8B-Instruct using the HotpotQA and MultiFieldQA datasets. Experimental results are shown in Table 5.
[…]